# QCD Axion Strings or Seeds?

### v1

@@ 1 Introduction
The QCD axion strings we are focussing on are actually one example among the possible objects that can play the role of impurities for cosmological phase transitions. The relevance of inhomogeneous, or seeded, transitions in this context has been in fact recognized long time ago Steinhardt:1981mm ; Steinhardt:1981ec ; Hosotani:1982ii ; Jensen:1982jv ; Witten:1984rs ; Yajnik:1986tg ; Yajnik:1986wq ; Preskill:1992ck . Topological defects, such as cosmic strings Yajnik:1986tg ; Yajnik:1986wq ; BhowmikDuari:1993np ; Dasgupta:1997kn ; Kumar:2008jb ; Koga:2019mee ; Duari:1994jh ; Lee:2013zca , domain walls Blasi:2022woz ; Blasi:2023rqi ; Agrawal:2023cgp , and monopoles Steinhardt:1981mm ; Steinhardt:1981ec ; Kumar:2009pr ; Kumar:2010mv ; Agrawal:2022hnf are among the most natural impurities arising from field theory. Compact objects, including (primordial) black holes, have been shown to affect the bubble nucleation rate as well Hiscock:1987hn ; Berezin:1987ea ; Arnold:1989cq ; Berezin:1990qs ; Metaxas:2000qf ; Pearce:2012jp ; Gregory:2013hja ; Burda:2015isa ; Mukaida:2017bgd ; Canko:2017ebb ; Kohri:2017ybt ; Oshita:2018ptr ; Balkin:2021zfd ; Shkerin:2021zbf ; Shkerin:2021rhy ; Briaud:2022few ; Jinno:2023vnr . Other possible sources are primordial density fluctuations Jinno:2021ury and energetic particle rays Kuznetsov:1997az ; Enqvist:1997wv ; Demidov:2015bua ; Strumia:2023awj .
@@ 2 Setup
We finally emphasize that the PQ–Higgs portal interaction is not protected by any symmetry and is in fact unavoidably generated by loops of the KSVZ fermions responsible for the mixed PQ–QCD anomaly. For instance, KSVZ fermions coupled to the PQ field with a yukawa interaction $y=M_{\psi}/f_{a}$ and to the SM only via QCD contribute to the portal by a three-loop diagram involving top quarks, which can be estimated as $\kappa_{\text{rad}}\sim 10^{-5}(M_{\psi}/f_{a})^{2}$. Larger contributions are expected if the KSVZ fermions are charged under the EW group. In addition, depending on the specific UV theory one generically expects other contributions to this coupling from physics at scales higher (or equal) than $f_{a}$. For this reason, will treat the portal coupling $\kappa$ (as well as the PQ self–quartic $\eta$) as a free parameter in our analysis.
@@ 4 Thermal history of SM plus PQ
Let us conclude this section by noticing that the spontaneous breakdown of the EW symmetry at the core of string C at temperatures much above the EW scale can potentially modify the QCD axion string dynamics in the early universe, e.g. in terms of friction with the thermal plasma, or other properties of the network such as particle production. A study of these features is left for future work.
@@ 6 Conclusion
Let us finally mention that while we have restricted our study to KSVZ–like models where the Higgs is neutral under the PQ symmetry, we expect similar implications for the electroweak phase transition also in DFSZ–like models where the Higgs doublets have additional couplings with the string due to the non–zero PQ charge.

### v2

@@ 1 Introduction
The QCD axion strings we are focussing on are actually one example among the possible objects that can play the role of impurities for cosmological phase transitions. The relevance of inhomogeneous, or seeded, transitions in this context has been in fact recognized long time ago Steinhardt:1981mm ; Steinhardt:1981ec ; Hosotani:1982ii ; Jensen:1982jv ; Witten:1984rs ; Yajnik:1986tg ; Yajnik:1986wq ; Preskill:1992ck . Topological defects, such as cosmic strings Yajnik:1986tg ; Yajnik:1986wq ; BhowmikDuari:1993np ; Dasgupta:1997kn ; Kumar:2008jb ; Koga:2019mee ; Duari:1994jh ; Lee:2013zca , domain walls Blasi:2022woz ; Blasi:2023rqi ; Agrawal:2023cgp , and monopoles Steinhardt:1981mm ; Steinhardt:1981ec ; Kumar:2009pr ; Kumar:2010mv ; Agrawal:2022hnf are among the most natural impurities arising from field theory. Compact objects, including (primordial) black holes, have been shown to affect the bubble nucleation rate as well Hiscock:1987hn ; Berezin:1987ea ; Arnold:1989cq ; Berezin:1990qs ; Metaxas:2000qf ; Pearce:2012jp ; Gregory:2013hja ; Burda:2015isa ; Mukaida:2017bgd ; Canko:2017ebb ; Kohri:2017ybt ; Oshita:2018ptr ; Balkin:2021zfd ; Shkerin:2021zbf ; Shkerin:2021rhy ; Strumia:2022jil ; Briaud:2022few ; Jinno:2023vnr . Other possible sources are primordial density fluctuations Jinno:2021ury and energetic particle rays Kuznetsov:1997az ; Enqvist:1997wv ; Demidov:2015bua ; Strumia:2023awj .
@@ 2 Setup
We finally emphasize that the PQ–Higgs portal interaction is not protected by any symmetry and is in fact unavoidably generated by loops of the KSVZ fermions responsible for the mixed PQ–QCD anomaly  [^4] Neglecting QCD as well as possible mixing between the KSVZ and the SM quarks, the portal is however technically natural Foot:2013hna .. For instance, KSVZ fermions coupled to the PQ field with a yukawa interaction $y=M_{\psi}/f_{a}$ and to the SM only via QCD contribute to the portal by a three-loop diagram involving top quarks, which can be estimated as $\kappa_{\text{rad}}\sim 10^{-5}(M_{\psi}/f_{a})^{2}$. Larger contributions are expected if the KSVZ fermions are charged under the EW group. In addition, depending on the specific UV theory one generically expects other contributions to this coupling from physics at scales higher (or equal) than $f_{a}$. Let us however note that as the effect of the axion strings on the electroweak sector will be controlled by the mass of the PQ radial mode, $m_{\rho}$, and by the ratio of couplings $\kappa/\eta$, a small portal can still be relevant depending on the size of the PQ self interaction. For this reason, we will treat the portal coupling $\kappa$ (as well as the PQ self quartic $\eta$) as a free parameter in our analysis.
@@ 4 Thermal history of SM plus PQ
Let us conclude this section by noticing that the spontaneous breakdown of the EW symmetry at the core of string C at temperatures much above the EW scale can potentially modify the QCD axion string dynamics in the early universe, e.g. in terms of friction with the thermal plasma, or other properties of the network such as particle production. In fact, as the SM particles in the plasma can now efficiently scatter off the Higgs core inside the string, we can expect the network to be friction dominated from temperatures $T\sim T_{r}^{\rm B}$ until $T\sim\sqrt{G\mu}f_{a}$ilenkin:2000jqa , where $\mu$ is the tension of string C (expected to be the same as the standard axion string up to small corrections in our setup) and $G$ is Newton’s constant. Depending on the value of $f_{a}$ the network may still be friction dominated (or anyway far from the standard scaling solution) at the time of axion domain–wall formation around the QCD crossover, thus modifying its energy density and therefore the abundance of emitted axions. In addition, the presence of a non–trivial Higgs profile in association with the axion–Higgs portal may provide an additional channel for producing axion excitations.
@@ 6 Conclusion
Let us also notice here that the presence of the portal interaction between the Higgs and the PQ scalar may actually modify the scale where the Higgs quartic turns negative due to the running in the pure SM Buttazzo:2013uya because of additional contributions and threshold corrections that improve the stability of the SM vacuum Elias-Miro:2012eoi . This provides motivation for considering a scenario with a sizable portal interaction  [^10] We thank the Referee on this point..
@@
Finally, let us mention that while we have restricted our study to KSVZ–like models where the Higgs is neutral under the PQ symmetry, we expect similar implications for the electroweak phase transition also in DFSZ–like models where the Higgs doublets have additional couplings with the string due to the non–zero PQ charge. This may result in more operators that can be possibly included in the Higgs–string EFT in Eq. (26).